OpenStudy (lgbasallote):

what is the general form of an nth order linear differential equation with constant coefficient? my notes says it's \[\large (A_n D^n + A_{n-1} D^{n-1} + \cdots + A_n)y = R(x)\] but a quiz i secretly acquired says it's \[\large (A_o D^n + A_1 D^{n-1} + \cdots + A_{n-1} D + A_n)y = R(x)\] so which one is right?

OpenStudy (anonymous):

first one is An An-1 ......... A0 i think ha?

OpenStudy (lgbasallote):

heh?

OpenStudy (lgbasallote):

yeah i think..i may have copied it wrong

OpenStudy (anonymous):

anyway those are constants it does'nt matter how to place them (any order is correct) so both are correct

OpenStudy (lgbasallote):

really? so what will be the wrong equation? can you give a sample?

OpenStudy (anonymous):

It is a notational problem. Both notations are OK.

OpenStudy (lgbasallote):

i think it would be nice to see a situation wherein it's wrong...so i;ll know when it's right..when it's wrong..

OpenStudy (anonymous):

there is no wrong situation...lol

OpenStudy (lgbasallote):

really? so it's always right?

OpenStudy (lgbasallote):

lol there must be a wrong situation :/

OpenStudy (anonymous):

Here are two representation of a general polynomial of degree 10 \[ p(x)=a(10) x^{10}+a(9) x^9+a(8) x^8+a(7) x^7+a(6) x^6+a(5) x^5+\\a(4) x^4+a(3) x^3+a(2) x^2+a(1) x+a(0)\] \[ q(x)= a(0) x^{10}+a(1) x^9+a(2) x^8+a(3) x^7+a(4) x^6+\\a(5) x^5+a(6) x^4+a(7) x^3+a(8) x^2+a(9) x+a(10) \]